### v2

# Glimpse: Generalized Local Imaging with MLPs

Abstract
Deep learning has become the state-of-the-art approach to medical tomographic imaging. A common approach is to feed the result of a simple inversion, for example the backprojection, to a (multiscale) convolutional neural network (CNN), which then computes the final reconstruction. Despite strong results on in-distribution test data similar to the training data, they overfit certain large-scale structures which leads to poor generalization on out-of-distribution (OOD) samples. Moreover, their memory complexity and training time scale unfavorably with image resolution, making them impractical for application at realistic clinical resolutions, especially in 3D. A standard U-Net requires a substantial 140GB of memory and 2600 seconds per epoch on a research-grade GPU when training on 1024 $\times$ 1024 images with batch size 64. In this paper, we introduce Glimpse, a local processing neural network for computed tomography which reconstructs a pixel value by processing only the measurements associated with the neighborhood of the pixel with a simple multi-layer perception (MLP). While achieving performance comparable to or better than successful CNNs like U-Net on in-distribution test data, Glimpse significantly outperforms them on OOD samples while maintaining a memory footprint almost independent of image resolution; 5GB memory suffices to train on 1024 $\times$ 1024 images, with each epoch requires 420 seconds. Because we built Glimpse to be fully differentiable it can also be used as a plug-in component of arbitrary deep learning architectures, enabling feats such as correction of miscalibrated projection orientations.
Index Terms:
Deep Learning, Computed Tomography, MLP, Uncalibrated Imaging
I Introduction
Convolutional neural networks (CNNs) have become the standard approach for tomographic image reconstruction [1]. U-Net [2] has emerged as an architecture underpinning numerous deep learning reconstruction methods, applied with great success to a variety of imaging problems including computed tomography (CT) [3], magnetic resonance imaging (MRI) [4] and photoacoustic tomography [5]. Its success is often attributed to its particular multi-scale architecture [6].
At the same time, certain aspects of multi-scale CNNs complicate their application to real problems. Despite good performance on in-distribution test images similar to the training data, they often overfit specific image content resulting in poor generalization to distribution shifts in image content and sensing as shown in this paper. Model-based networks attempt to address this drawback by integrating the forward and adjoint operators into multiple network layers or iterations [7, 8, 9, 10, 11, 12]. However, the required memory for CNNs directly scales with image resolution [13]. For instance, the widely used U-Net requires a substantial 140GB memory and 2600 seconds per epoch when training on 1024 $\times$ 1024 images using two Tesla A100 GPUs. This latter drawback is further exacerbated with model-based networks such as learned primal-dual (LPD) [8], which achieves strong performance but requires over 80GB memory and very long training time even at a lower resolution of $512\times 512$. This increased memory demand is due to the repeated application of the forward model and its adjoint in forward and backward passes of the neural network. This makes standard CNN-based pipelines impractical for real-world scenarios involving resolutions beyond $512\times 512$.
Figure 1: A point source image, its sinogram, and the sparse view FBP reconstruction. While the corresponding measurements for this pixel have sinusoidal support in the sinogram, this information is diffused all over the FBP image. The contrast of the FBP image has been stretched to emphasize this effect.
To better understand the mechanics behind the poor generalization of U-Net-like CNNs which compute the reconstruction from filtered backprojections (FBP) [14], we designed an experiment as follows. Figure 1 shows an object with a point source, its sparse view sinogram measurements with sinusoidal support, and the FBP reconstruction. It is evident that the FBP is supported over the entire field of view. This observation raises the question of the ideal receptive field size for CNNs like U-Net: a large receptive field may be beneficial to capture all information correlated with the value of a target pixel [15, 16].
However, models with large receptive fields often overfit specific image content in training data which leads to poor generalization on out-of-distribution samples [17]. Indeed, Figure 2 shows that while U-Net produces good results when tested on in-distribution data similar to training data (here chest images), it performs poorly on out-of-distribution (here brain images). This makes CNNs like U-Net problematic in domains such as medical imaging where robustness over distribution shifts and other uncertain and variable factors is of relevance [18].
Figure 2: Performance of U-Net [2] trained on chest images: evaluation on in-distribution test data (chest samples) and OOD brain samples shows that the large receptive field of U-Net hinders its ability to generalize on OOD samples, with its PSNR even falling below that of FBP reconstruction.
In this paper, we introduce a new deep learning imaging architecture termed Glimpse—a simple local processing neural network adapted to the geometry of computed tomography. As shown in Figure 3, to recover the image intensity at a given target pixel, we use a MLP that takes only the local sinogram measurements associated with this pixel and its neighbors. There is no backprojection step. This localization results in robust performance, particularly when dealing with OOD data.
Figure 3: Glimpse; a single MLP processes the measurements associated with the pixel $(x,y)$ and its neighbors extracted from the sinogram. This local processing network has promising performance on OOD data while being computationally efficient all due to its locality.
At the same time, this design makes Glimpse highly computationally efficient; it permits training on mini-batches of both pixels and objects. This flexibility leads to fast and efficient training, requiring a small, fixed amount of memory almost independent from the image resolution. This allows training Glimpse on large, realistic images in resolution $1024\times 1024$ and beyond.
We built Glimpse to be fully differentiable, all the way down to the sensing and integration geometry. This has several advantages over the standard CNN-based architectures. For instance, most methods for CT image recovery strongly rely on the sensor geometry information encoded in the forward operator, whether explicitly, as seen in methods like FBP [14], SART [19], LGS [7], and LPD [8] or implicitly as used in U-Net [2] when taking FBP as input. This fixed geometry is a problem when faced with uncertainties in calibration or blind inversion problems where the sensor geometry information is entirely unavailable. While such uncertainties might degrade the quality of reconstructions of the standard methods [20, 21], our differentiable architecture allows the optimization of projection angles which can estimate the right projection angles and improve the quality of reconstructions.
II Related Work
Model-based vs Model-free Inversion. There are two major classes of deep-learning-based approaches to CT: model-based and model-free inversion. In the model-based approach, neural networks process raw sinograms and map them to the desired CT images while the Radon transform is integrated into multiple network layers or iterations [7, 8, 11]. These methods perform remarkably well across various inverse problems, but they are computationally expensive, especially during training [13]. The high computational cost is due, among other factors, to the repeated application of the Radon transform and its adjoint in forward and backward passes of the neural networks.
In contrast, model-free approaches offer a computationally cheaper alternative. The Radon transform (or its adjoint) is only used once in FBP computation before the neural network [3, 22, 23]. However, these models often require deep networks with a large receptive field to leverage the information delocalized across the FBP image. Recently, Hamoud et al. [16] used a measurement rearrangement technique to stratify backprojected features by angle and thus enable the use of smaller, shallower CNNs.
MLPs for Imaging A multi-layer perceptron (MLP) is a fundamental neural network architecture used in a great variety of applications. Recently, vision transformers [24] and MLP-mixers [25] have shown promising performance in various computer vision tasks like image classification [26] and image restoration [27]. While, unlike CNNs, a vanilla MLP lacks a good inductive bias for imaging, in particular translation equivariance, vision transformers and MLP-mixers restore it by processing patches instead of entire images [28]. However, these strategies require large datasets and networks to achieve performance comparable with CNNs. In our work, we propose a differentiable local processing network for CT imaging, demonstrating that even a small MLP can achieve performance on par with or even surpassing that of popular image-to-image CNNs.
Uncalibrated CT Imaging. In CT imaging, the acquisition operator is often known but an insufficient number of measurements is obtained. This may occur when a reduced number of projections is used to minimize radiation exposure or shorten acquisition time (sparse view) or when only a limited cone of projection angles may be used (limited view). In certain situations, the acquisition operator is only partially or approximately known. Neglecting this uncertainty can result in a significant drop in the quality of the reconstructions [20]. To tackle this challenge, total least squares approaches have been developed, involving the perturbation of an assumed forward operator [29, 30, 31]. Recently, Gupta et al. [21] used autodifferentiation and gradient descent to estimate the uncalibrated forward operator in a self-consistent manner.
III Computed Tomography
CT imaging [32] plays an important role in many applications including medical diagnosis [33], industrial testing [34], and security [35]. We consider 2D computed tomography where the image of interest $f({\mathbf{x}})$ with size $D\times D$ is reconstructed from measurements of (X-ray) attenuation. The forward model is the Radon transform $Rf$ which computes integrals of $f({\mathbf{x}})$ along lines $L$,
$Rf(L)=\int_{L}f\big{(}{\mathbf{x}}\big{)}|d{\mathbf{x}}|.$ (1)
We parameterize a line $L$ by its distance from the origin $t$ and its normal vector’s angle with the $x$-axis $\alpha$. We can then reformulate (1) as
[…]
where,
$x(z)$ $=z\cos(\alpha)-t\sin(\alpha),$ (3)
$y(z)$ $=z\sin(\alpha)+t\cos(\alpha).$ (4)
The image of interest is observed from a finite set of $r$ different viewing directions $\{\alpha_{m}\}_{m=1}^{r}$, each having $N$ parallel, equispaced rays. The measurements of the attenuation are then represented as a transform-domain “image” ${\mathbf{s}}\in\mathbb{R}^{N\times r}$ called a sinogram.
Standard methods for CT image recovery discretize the image of interest $f({\mathbf{x}})$ into a discrete image ${\mathbf{f}}\in\mathbb{R}^{N\times N}$ supported on an $N\times N$ grid. After discretization, the forward model can be written as
${\mathbf{s}}={\mathbf{A}}{\mathbf{f}}+{\mathbf{n}}.$ (5)
where ${\mathbf{A}}$ is the matrix of the discretized Radon transform and we model the measurement noise by ${\mathbf{n}}$. The most commonly used analytical inversion method is the filtered backprojection (FBP),
${\mathbf{f}}^{\text{FBP}}_{x,y}=\sum_{m=1}^{r}\tilde{{\mathbf{s}}}[y\cos(\alpha_{m})-x\sin(\alpha_{m}),m],$ (6)
where ${\mathbf{f}}^{\text{FBP}}\in\mathbb{R}^{N\times N}$ is the FBP reconstruction, $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$, ${\mathbf{h}}$ is a certain high-pass filter, * denotes the convolution and linear interpolation is used in (6) for evaluating $\tilde{{\mathbf{s}}}[x,\cdot]$ when $x$ is not an integer. As shown in Proposition 2 in Appendix -C, while the Ram–Lak filter is the optimal choice for ${\mathbf{h}}$ in the case of noise-free complete measurements, it can amplify the noise in real-world noisy measurements, leading to poor reconstruction.
With measurement noise and an incomplete collection of projections, tomographic image reconstruction from a sinogram becomes an ill-posed inverse problem that requires an image prior. In the following section, we introduce our proposed method, Glimpse, designed so that it respects the geometry of CT imaging.
III-A Glimpse: Local Imaging with MLPs
To recover the image ${\mathbf{f}}(x,y)$ at location ${\mathbf{x}}=(x,y)$, we identify the elements in the sinogram ${\mathbf{s}}$ influenced by this pixel. As illustrated in Figure 1, the corresponding measurements for the pixel $(x,y)$ are supported along a sinusoidal curve in the sinogram; we denote them $\text{SIN}_{x,y}\in\mathbb{R}^{r}$, with elements being given as
$\text{SIN}_{x,y}[m]={\mathbf{s}}[y\cos(\alpha_{m})-x\sin(\alpha_{m}),m].$ (7)
Similar to (6), we can use interpolation to evaluate ${\mathbf{s}}[x,\cdot]$ for non-integer $x$. This localization is formally captured by the following proposition.
Proposition 1 (Impulse response of Radon transform).
Let $f(u,v)=\delta(u-x,v-y)$ be the Dirac delta distribution in $\mathbb{R}^{2}$ at location $(x,y)$. Its Radon transform (in the sense of distributions) is
$Rf(\alpha,t)$ =$=\begin{cases}1,&\text{if}\ t=r\cos(\alpha+\varphi)\\ 0,&\text{otherwise},\end{cases}$
[…]
The proof is standard and outlined for completeness in Appendix -D.
The sinusoidal portion of the sinogram $\text{SIN}_{x,y}$ should have enough information to recover the pixel intensity $(x,y)$ as it contains all the measurements associated with this pixel. Note however that the pixel at $(x,y)$ influences the integral over any line passing through it and thus also the parts of the sinogram corresponding to pixels on those lines. This can be loosely thought of as a consequence of non-orthogonality of the Radon transform. The above statement is thus more precisely a statement about the filtered sinogram since information is “relocalized” by the high-pass filtering step in the FBP.
This is related to the celebrated support theorems of Sigurdur Helgason, Jan Boman, and others [36, 37, 38, 39]. These theorems state that under appropriate conditions a compactly-supported image may be recovered from a compactly-supported subset of its Radon data. These results do not involve filtering explicitly, but its influence is implicit. They apply to idealized sampling and SNR conditions.
Indeed, the high-pass filtering in the FBP is derived for noiseless data and a continuum of observed angles. In reality the projections are corrupted with noise and come from a sparse subset of projection angles. We address this by 1) incorporating “contextual information” about the target pixel and 2) letting the filter be learnable to adapt it to the specifics of discretization and noise.
As shown in Figure 3, we exploit the spatial regularity of medical images (encoded in training data) by using the measurements which provide local information around $(x,y)$. This ensures that the model does not overfit large-scale features in the training data while maintaining low computational complexity. We thus additionally extract from the sinogram the regions associated with the neighboring pixels around $(x,y)$ and store this information in vector ${\mathbf{p}}_{x,y}$,
${\mathbf{p}}_{x,y}=\{\text{SIN}_{x+dn,y+dn^{\prime}}|n,n^{\prime}=-\left\lfloor C/2\right\rfloor,\cdots,\left\lfloor C/2\right\rfloor\},$ (8)
where $K=C^{2}$ determines the number of neighboring pixels around $(x,y)$ for an odd number $C\geq 1$ and $d$ denotes the scale of the window which adjusts the receptive field. In order to recover the image at pixel $(x,y)$ from ${\mathbf{p}}_{x,y}$, we use a multi-layer perception $\text{MLP}_{\theta}:~{}\mathbb{R}^{r\times K}\to~{}\mathbb{R}$ parameterized by $\theta$,
$\hat{{\mathbf{f}}}(x,y)=\text{MLP}_{\mathrm{\bm{\theta}}}\big{[}{\mathbf{p}}_{x,y}\big{]},$ (9)
which estimates the pixel intensity $\hat{{\mathbf{f}}}_{x,y}$ from the local features around $(x,y)$. We call the proposed model Glimpse, standing for generalized [^1] The word “generalized” emphasizes that locality is also encoded in the transform domain, not just in real space as in some of earlier work. local imaging with MLPs. In the following section, we describe how our implementation of Glimpse allows to adapt to noisy measurements. We then propose a training strategy with resolution-agnostic memory usage in Section III-C. In Appendix -B, we show how Glimpse compensates for calibration errors. Further details for network architecture and training can be found in Appendix -A.
III-B Adaptive Filtering for Noisy Measurements
The Ram–Lak high-pass filter is the optimal filter ${\mathbf{h}}$ for the FBP reconstruction in the case of complete noise-free measurements; see Appendix -C for a standard demonstration. In real applications, however, we always encounter noisy projections from a subset of angles. The Ram–Lak filter is then suboptimal and typically degrades the reconstruction quality as it amplifies high-frequency noise. Alternative filters with lower amplitudes in high frequencies like Shepp–Logan, cosine, and Hamming have been used to mitigate the noisy measurements, but they are all ad hoc choices. It is advantageous to adapt ${\mathbf{h}}$ to the specifics of noise and sampling strategy in the target application. To design this task-specific filter, we let $\text{MLP}_{\mathrm{\bm{\theta}}}$ take as input the filtered sinogram $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$ and consider the filter ${\mathbf{h}}$ (in Fourier space) as trainable parameters to be optimized during training. This allows us to automatically learn a noise-adaptive filter from data, again with almost no additional computational cost.
III-C Resolution-agnostic Memory Usage in Training
To simplify notation, we denote the entire Glimpse pipeline described above by $\hat{{\mathbf{f}}}({\mathbf{x}})=\text{{Glimpse}}_{\phi}({\mathbf{x}},{\mathbf{s}})$. The inputs are the target pixel coordinates ${\mathbf{x}}=(x,y)$ and the sinogram ${\mathbf{s}}$; the output is an estimate of ${\mathbf{f}}(x,y)$. The parameters $\phi$ denote the trainable parameters of Glimpse including the MLP weights $\theta$, the projection angles $\{\alpha_{m}\}_{m=1}^{r}$ (see Appendix -B), the adaptive filter ${\mathbf{h}}$ and the window receptive field scale $d$. We consider a set of training data $\{({\mathbf{s}}_{i},{\mathbf{f}}_{i})\}_{i=1}^{L}$ from the noisy sinograms and images. We optimize the Glimpse parameters $\phi$ using gradient-based optimization to minimize
$\phi^{*}=\operatorname*{argmin}_{\phi}~{}\sum_{i=1}^{N^{2}}\sum_{j=1}^{L}|\text{{Glimpse}}_{\phi}({\mathbf{x}}_{i},{\mathbf{s}}_{j})-{\mathbf{f}}_{j}({\mathbf{x}}_{i})|^{2}.$ (10)
At inference time, we simply evaluate the image intensity at any pixel as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{{Glimpse}}_{\phi^{*}}({\mathbf{x}},{\mathbf{s}}_{\text{test}})$. One major advantage of Glimpse compared to CNNs like U-Net and LPD is its memory and compute complexity. CNN-based models exhibit memory requirements that scale directly with image resolution, making them prohibitively expensive for realistic image resolutions. As shown in (10), Glimpse can be trained using stochastic gradient-based optimizers with the flexibility to select mini-batches from both the objects and pixels. This adaptability in mini-batch pixel selection grants a memory footprint agnostic to resolution making Glimpse suitable for training on realistic image resolutions like $1024\times 1024$ and higher.
IV Experiments
We simulate parallel-beam X-ray CT with $r=30$ projections uniformly distributed around the object with additive Gaussian noise to reach a signal-to-noise ratio (SNR) of 30 dB. The reconstruction quality is quantified using the peak signal-to-noise ratio (PSNR) and Structural Similarity Index (SSIM) [40]. We compare the performance of Glimpse with successful CNN-based models: U-Net [2], learned gradient scheme (LGS) [7] and learned primal-dual (LPD) [8] for sparse view CT image reconstruction. We use 35820 training samples of chest images from the LoDoPaB-CT dataset [41] in resolution $128\times 128$. Model performance is assessed on 64 in-distribution test samples of chest images, while 16 OOD brain images [42] are included to evaluate the generalization capability of the models. For further information regarding the network architectures and training details please refer to Section -A.
In Section IV-A, we compare Glimpse to CNN-based models for sparse view CT image reconstruction on both in-distribution and OOD data. In Section IV-B, we analyze the computational efficiency of the aforementioned models. We analyze the learned filters ${\mathbf{h}}$ across different measurement noise levels in Section IV-C. In Appendix -B we consider the uncalibrated and blind scenarios.
IV-A Sparse view CT Image Reconstruction
Figure 4: Performance of different models trained on training data of chest images evaluated on in-distribution and OOD samples for sparse view CT image reconstruction. Glimpse has excellent performance on OOD data due to its localized MLP, significantly better than U-Net [2] and LGS [7] and comparable with LPD [8].
TABLE I: Comparison of different models for sparse view CT image reconstruction
(a) The reconstruction quality averaged on 64 test samples
In-distribution (chest) Out-of-distribution (brain)
PSNR SSIM PSNR SSIM
FBP [14] 17.0 0.17 17.1 0.22
U-Net [2] 30.1 0.84 15.1 0.28
LGS [7] 30.9 0.84 20.5 0.54
LPD [8] 31.6 0.86 25.5 0.76
Glimpse 30.9 0.84 25.1 0.79
(b) Memory usage and training time (batch size 64)
Glimpse U-Net [2] LGS [7] LPD [8]
Num params 900k 7800k 19k 400k
$128\times 128$ 4GB / 114s 6GB / 34s 4GB / 384s 13GB / 963s
$256\times 256$ 4GB / 123s 16GB / 117s 13GB / 575s 41GB / 1517s
$512\times 512$ 4GB / 185s 53GB / 460s 45GB / 1682s $>80$GB
$1024\times 1024$ 5GB / 419s $>80$GB $>80$GB $>80$GB
The upper row of Figure 4 and Table I(a) show the performance of different models on in-distribution test samples of chest images. This experiment shows that Glimpse, by leveraging only a single MLP network, can outperform successful CNNs like U-Net and achieve comparable performance with LGS and LPD methods.
The lower row of Figure 4 and Table I(a) shows a comparison of the performance of various models trained on chest images when applied to OOD brain images. This experiment demonstrates that while U-Net excels on in-distribution samples, its performance significantly deteriorates on OOD data.
On the contrary, Glimpse shows strong performance on OOD data. Although LPD’s performance on OOD data is sometimes comparable or slightly better than that of Glimpse, it comes at a very high memory and compute cost due to the repeated application of the forward operator and its adjoint in the network architecture; we analyze this in the next section.
IV-B Computational Efficiency
The fact that LPD far outperforms U-Net on OOD data is a testament to the benefits of incorporating the forward operator in the architecture. On the other hand, as evident from Table I(b), it comes at the cost of unfavorable training time and memory footprint which rapidly worsens with resolution. Table I(b) shows that CNN-based models may become impractical already at resolutions like $512\times 512$, even on GPUs with 80GB memory.
On the other hand, Glimpse is computationally efficient; the memory usage remains almost unaffected by image resolution. Remarkably, Glimpse can be trained with only 5GB memory in less than a day, even when dealing with resolutions of $1024\times 1024$ and higher. Figure 5 shows the performance of Glimpse on in-distribution and OOD samples in resolution $512\times 512$ where we considered 40dB measurement noise. This experiment demonstrates that a relatively small MLP, with almost 10 times fewer parameters than a standard U-Net, can achieve strong performance in realistic high resolutions while maintaining a rather modest memory footprint.
(a) In-distribution chest samples
(b) OOD brain samples
Figure 5: Glimpse’s performance in resolution $512\times 512$ trained on chest training data with $r=30$ projections and 40dB noise; Glimpse requires only 4GB memory and can be trained in less than 10 hours on a single GPU.
IV-C Learned Filter
Figure 6: The learned filter for datasets with different noise levels, all the filtered are initialized by Ram-Lak filter in Glimpse architecture. By increasing the noise level, the filter assigns smaller amplitudes for high-frequencies to suppress the noise and aligns with the optimality of the Ram-Lak filter for noise-free complete measurements shown in Section -C.
In this section, we analyze the behavior of the learned filter obtained through training of Glimpse introduced in section III-B across datasets with different measurement noise levels. This analysis provides useful signal processing insights into how the properties of the learned filter are influenced by varying noise levels.
In Figure 6 we show the frequency response of the learned filters, alongside with standard hand-crafted filters such as Ram-Lak, Shepp-Logan, and Hamming filters. These learned filters are derived from Glimpse training on datasets characterized by different levels of noise.
As expected from the discussion in Appendix -C, the learned filter for noise-free measurements is similar to the Ram–Lak filter, with a relatively high amplitude in high frequencies. As the noise level increases (by decreasing the noise SNR), the filter progressively takes smaller values in high frequencies to suppress the noise. This confirms that Glimpse can autonomously adapt the characteristics of the filter according to the noise level observed in the training data.
V Limitations and Conclusion
We used a natural notion of locality for CT which is adapted to sinogram geometry. This is different from CNNs that reconstruct the image as a whole, and where the notion of locality (at small scales) is in the sense of real image space. Our approach adopts a coordinate-based strategy, focusing on processing the sinusoidally-shaped regions of the sinogram associated with pixels using a small MLP. Our results demonstrate that this localized processing framework does significantly improve robustness to OOD data while maintaining nearly constant memory requirements across different resolutions, being computationally efficient even at realistic high resolutions. The differentiable architecture makes it easy to combine Glimpse with other deep learning pipelines. Concretely, we demonstrated the possibility to learn sensor geometry for uncalibrated systems and to adapt the learned filter according to the noise level and projection sparsity.
While the memory required by Glimpse varies little with image resolution, a drawback of the current scheme is that memory and computing costs increase as the number of projections $r$. A possible alternative to the standard MLP architecture which is the culprit for this is to use mixture-of-experts layers [43, 44, 45], which selectively employ smaller MLPs for processing inputs. This mixture-of-experts approach is an effective drop-in replacement for standard MLP layers of language transformers [46] and vision transformers [24].
Glimpse could be integrated with various imaging problems involving line integrals in forward operators such as photoacoustic [47, 48] and cryo-electron tomography (cryoET) [49, 50]. Its future full-3D adaptation may yield efficient architectures which resolve the fundamental memory issues with applications of deep learning in 3D medical imaging. This extension is particularly interesting given the ability of Glimpse to operate locally and its near-fixed memory requirement across resolution, which makes it an ideal choice for large 3D objects.
-A Network Architecture and Training Details
For Glimpse architecture, we use an MLP network comprising 9 hidden layers, each with dimensions [256, 256, 256, 256, 128, 128, 128, 64, 64] with ReLu activations. The input to the MLP network consists of sinusoidal curves sampled from $K=9^{2}$ neighboring pixels. To prevent edge artifact of the circular convolution, we apply zero-padding with a size of 512 to the sinogram before applying the filter ${\mathbf{h}}$. Linear interpolation is used in (7). For the experiment in resolution $512\times 512$ in Section IV-B, we use a larger network with hidden layer dimensions [1024, 1024, 1024, 512, 512, 512, 256, 256] to enhance the quality of reconstructions.
We implement our model in PyTorch [51] on a machine equipped with a Nvidia A100 GPU with 80GB of memory to train the different architectures. We report the maximum capacity of the graphics card during training and the time needed to complete the training. All models in Section IV were trained for 200 epochs with MSE loss using the Adam optimizer [52]. A learning rate $10^{-4}$ was used for Glimpse and U-Net, while LGS and LPD were trained with a learning rate $10^{-3}$. All models were trained with batch size 64. In the case of Glimpse, for each mini-batch of random objects, we performed optimization on a random mini-batch of 512 pixels 3 times.
-B Learned Sensor Geometry
CT imaging algorithms such as FBP [14], SART [19], LGS [7], LPD [8] assume that the projection angles $\{\alpha_{m}\}_{m=1}^{r}$ are known. In an uncalibrated system where sensor geometry is different from what the algorithms assume, the quality of reconstruction deteriorates [20, 53]. Glimpse allows directly optimizing the projection angles during training. We thus jointly optimize $\{\alpha_{m}\}_{m=1}^{r}$ with other trainable parameters in (10). This additional angle estimation incurs a very modest computational cost.
In the absence of calibration, we cannot expect to have paired ground truth images. In the following experiments, we only want to showcase the possibility to differentiably optimize over angles in Glimpse so we assume having access to paired data (while simulating the uncalibrated forward operator). In practice we could use a self-supervised loss, for example, based on equivariance [54].
(a) Given projection angles
(b) Estimated projection angles
(c) Reconstructions
Figure 7: Estimated sensor geometry by Glimpse (LSG) and reconstructions for an uncalibrated system with a random sensor shift; as expected, the learnable sensor geometry can effectively learn the projection angles and exhibits excellent robustness with no degradation under such a big model mismatch and measurement noise (30dB).
We assess the performance of Glimpse in situations with mismatched projection orientations. In the following experiments, we place $r=30$ sensors uniformly around the object at angles $\alpha=0^{\circ},6^{\circ},...,174^{\circ}$. We conduct a comparative analysis of three models: 1) Glimpse (vanilla), with no learnable sensor geometry, 2) Glimpse (LSG), incorporating the proposed learned sensor geometry, and 3) Glimpse (calibrated), operating under ideal conditions with no model mismatch (informed with correct projection angles). We let the Glimpse (LSG) learn the projection angles from the training data where the optimized values $\{\alpha_{m}\}_{i=1}^{r}$ obtained through training can provide a reliable estimate of the actual projection angles.
Uncalibrated system with random sensor shifts: As shown in Figure 7(a), we randomly perturb projection angles by a normally distributed error so that $\alpha_{i}^{\text{given}}=\mathcal{N}(\alpha_{i},\sigma^{2})$; we set $\sigma=2^{\circ}$. We initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with $\alpha_{i}^{\text{given}}$. Figure 7(b) shows the estimated projection angles obtained through training—Glimpse (LSG) accurately recovers the angles even in the presence of 30 dB measurement noise. As shown in Figure 7(c), this accurate estimation of projection angles results in high-quality reconstructions by Glimpse (LSG) comparable with the network trained in an ideal calibrated system.
Blind inversion with no information from projection angles: We consider the blind scenario where the model operates without any prior knowledge of the sensor geometry making inversion challenging. As shown in Figure 8(a), we initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with random values. The estimated projection angles are shown in Figure 8(b), highlighting Glimpse (LSG)’s ability for data-driven sensor geometry estimation. Figure 8(c) presents the reconstructions achieved by Glimpse in both its vanilla and LSG versions. As expected, FBP and the Glimpse (vanilla) show poor reconstructions due to the missing sensor geometry information. On the other hand, Glimpse (LSG) could accurately reconstruct both in-distribution and OOD samples. Remarkably, these results are comparable to those achieved by the calibrated Glimpse with informed projection angles.
(a) Given projection angles
(b) Estimated projection angles
(c) High-quality reconstructions by Glimpse (LSG) despite having no information from sensor geometry.
Figure 8: Estimated sensor geometry by Glimpse (LSG) and reconstructions for blind inversion; Glimpse (LSG) was initialized with random projection angles $\{\alpha_{m}\}_{i=1}^{r}$ (a) could reliably estimate the projection angles purely from data (b) resulting in high-quality reconstructions (c).
-C Optimal Filter for FBP Reconstruction
Proposition 2 (Reconstruction for continuous Radon transform).
We have the following identity
[…]
where $\psi$ is the filter that has for Fourier transform ||$\cdot$|| ⋅ |.
Proof.
Let ${\mathbf{p}}=(x,y)$, $\mathrm{\bm{\xi}}=(\xi_{1},\xi_{2})$. We have
[…]
$=$ $\int_{0}^{+\infty}\int_{0}^{2\pi}\mathcal{F}_{2D}(f)(r\cos(\theta),r\sin(\theta))$
$\exp(2i\pi r\langle{\mathbf{k}},{\mathbf{p}}\rangle)rdrd\theta,$
by doing a change of variable in polar coordinates, where ${\mathbf{k}}=(\cos(\theta),\sin(\theta)).$ Observe that $\mathcal{F}_{2D}(f)(r\cos(\theta),r\sin(\theta))$ is the Fourier Transform of $f$ along the line of direction ${\mathbf{k}}$. By the Fourier slice theorem [32], we have
$\mathcal{F}_{2D}(f)(r\cos(\theta),r\sin(\theta))=\mathcal{F}_{1D}(Rf(\theta,\cdot))(r)$
By symmetry of the Radon transform, we have $Rf(\theta,r)=Rf(\theta+\pi,-r)$. Finally,
$f(x,y)=$ $\int_{-\infty}^{+\infty}\int_{0}^{\pi}\mathcal{F}_{1D}(Rf(\theta,\cdot))(r)\exp(2i\pi r\langle{\mathbf{k}},{\mathbf{p}}\rangle)$
$|r|drd\theta=\int_{0}^{\pi}\mathcal{F}_{1D}^{-1}\left(\mathcal{F}_{1D}(Rf(\theta,\cdot))\odot|\cdot|\right)d\theta.$
This shows that
[…]
where $\psi$ is the filter that has for Fourier transform ||$\cdot$|| ⋅ |.
-D Proof of Proposition 1
Proof.
Using the definition of the radon transform in (2), we have
$Rf(\alpha,t)=\int_{-\infty}^{+\infty}\delta($ $z\cos(\alpha)-t\sin(\alpha)-x,$
$z\sin(\alpha)+t\cos(\alpha)-y)dz.$
Solving $z\cos(\alpha)-t\sin(\alpha)-x=0$ for $z$ leads to
$z=\frac{t\sin(\alpha)+x}{\cos(\alpha)}.$
Then, solving $z\sin(\alpha)+t\cos(\alpha)-y=0$ for $t$, using the previous expression for $z$ leads to
$t=y\cos(\alpha)-x\sin(\alpha).$